### v1

# Machine Unlearning: A Comprehensive Survey

Abstract.
As the right to be forgotten has been legislated worldwide, many studies attempt to design unlearning mechanisms to protect users’ privacy when they want to leave machine learning service platforms. Specifically, machine unlearning is to make a trained model to remove the contribution of an erased subset of the training dataset. This survey aims to systematically classify a wide range of machine unlearning and discuss their differences, connections and open problems. We categorize current unlearning methods into four scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in unlearning. Since centralized unlearning is the primary domain, we use two parts to introduce: firstly, we classify centralized unlearning into exact unlearning and approximate unlearning; secondly, we offer a detailed introduction to the techniques of these methods. Besides the centralized unlearning, we notice some studies about distributed and irregular data unlearning and introduce federated unlearning and graph unlearning as the two representative directions. After introducing unlearning methods, we review studies about unlearning verification. Moreover, we consider the privacy and security issues essential in machine unlearning and organize the latest related literature. Finally, we discuss the challenges of various unlearning scenarios and address the potential research directions.
machine unlearning
ccs: Computing methodolgies Machine learningccs: Computing methodolgies Aritificial intelligenceccs: Security and privacyccs: Theory of computation
1. Introduction
Over the past decade, enormously increased data and fast hardware improvement have driven machine learning developments quickly. Machine learning (ML) algorithms and artificial intelligence (AI) are embedded into day-to-day applications and wearable devices (Mahdavinejad et al., 2018). It further collects more and more users’ information, including users’ private data, such as driving trajectory, medical and online shopping records. On the one hand, such an enormous amount of data helps to improve ML and AI development further. However, on the other hand, it threatens users’ privacy and demands the enormous requirement of data management to ensure information security and privacy in machine learning.
Machine unlearning is recently proposed to protect users’ privacy when they want their data forgotten by a trained machine learning model. It draws increasing research attention because Right to be forgotten was legislated recently in many countries, such as the General Data Protection Regulation (GDPR) (Mantelero, 2013) in the European Union, PIPEDA privacy legislation in Canada (of the Privacy Commissioner of Canada, 2018), and the California Consumer Privacy Act (de la Torre, 2018) in the United States. The legislation requires that companies must take reasonable measures to achieve the erasure of personal data concerning. For a machine learning application, it means that individual users have the right to ask companies to delete their private data once collected from them. The deletion is not only erasing their data from a database. It needs to remove the contribution of the erased subset of whole training data from the already trained model, which process was first conceptualized as machine unlearning (Cao and Yang, 2015). Specifically, if a user (Alice) wants to exercise her right to be forgotten (Mantelero, 2013) (e.g., quitting a ML application), then the trained model of such application must ”unlearn” her data. Such a process includes two steps: first, a subset of the whole dataset previously used for training a ML model is requested to be erased; second, the ML model provider removes the contribution of the erased data from the trained model. A naive data-erasing method is retraining a new model from scratch (Bourtoule et al., 2021). However, the computation and storage costs of retraining are expensive, especially in complex learning tasks.
Many researchers tried to find efficient and effective methods to realize unlearning rather than naive retraining, and there are several common challenges, which are summarized as follows. Stochasticity of training: A huge amount of randomness exists in the training process in machine learning, especially in complicated models’ training periods such as CNNs (Gu et al., 2018) and DNNs (Sze et al., 2017). This randomness makes the training result non-deterministic (Bourtoule et al., 2021) and raises difficulties for machine unlearning to estimate the contribution of the specific erased data sample. Incrementality of training: The training process in machine learning is an incremental procedure. In other words, the model update on the former data point will influence the contribution of the data points fed into the model after this data. Determining a way to effectively remove the influence of the to-be-erased data samples from the trained model is challenging for machine unlearning (Koh and Liang, 2017). Catastrophe of unlearning: Nguyen et al. (Nguyen et al., 2020) indicated that an unlearned model usually performs worse than the model retrained from scratch. Especially when an unlearning method tries to remove a huge amount of data samples, the degradation could be exponential. They referred to such sharp degradation as catastrophic unlearning (Nguyen et al., 2020). Although several studies have explored ways to mitigate unlearning degradation by bounding the loss function or restricting unlearning threshold, how to eliminate catastrophic unlearning is still an open problem. Recently, many studies put efforts into solving these three main challenges and proposed many novel mechanisms which promote the progress of machine unlearning.
[…]
ARC
Machine UnlearningCentralized UnlearningUnlearning CategoriesExact Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Approximate Unlearning: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021; Nguyen et al., 2020, 2022a; Fu et al., 2022; Khan and Swaroop, 2021) Unlearning TechniquesSplit Unlearning: (Cao and Yang, 2015; Bourtoule et al., 2021; Golatkar et al., 2021; Graves et al., 2021; Garg et al., 2020; Brophy and Lowd, 2021; Yan et al., 2022; Chen et al., 2022)Certified Data Removal: (Wu et al., 2020, 2022a; Guo et al., 2020; Sekhari et al., 2021; Neel et al., 2021; Schelter et al., 2021; Ginart et al., 2019; Golatkar et al., 2020a; Mehta et al., 2022; Li et al., 2021; Warnecke et al., 2021)Bayesian-based Unlearning: (Nguyen et al., 2020, 2022a; Fu et al., 2022)Distributed and Irregular Data UnlearningFederated Unlearning: (Wang et al., 2022; Liu et al., 2021, 2022a; Gupta et al., 2021; Wu et al., 2022b)Graph Unlearning: (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022)Unlearning Verification: (Hu et al., 2022; Thudi et al., 2022; Ma et al., 2022; Sommer et al., 2022; Huang et al., 2021a) Privacy and Security Issues in Machine UnlearningPrivacy Threats in UnlearningMembership Inference: (Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shokri et al., 2017; Lu et al., 2022; Huang et al., 2021a; Graves et al., 2021; Golatkar et al., 2020b)Privacy Reconstruction: (Gao et al., 2022; Chen et al., 2021b; Zanella-Béguelin et al., 2020; Shintre et al., 2019; Baumhauer et al., 2022; Graves et al., 2021; Golatkar et al., 2020b)Unlearning ApplicationsAnomaly Removal: (Wang et al., 2019; Liu et al., 2022b; Cao et al., 2018; Du et al., 2019; Shibata et al., 2021; Ye et al., 2022; Ganhör et al., 2022)Data Unlearnable: (Huang et al., 2021b)
Figure 1. Our taxonomy for machine unlearning. The introduction order will also follow this figure. We classify current unlearning literature into four main scenarios: centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
The goal of this survey is to classify and systematize machine unlearning according to the reseaching problems and goals in unlearning processes, and to review their differences and connections, advantages and disadvantages. We first catalogue unlearning methods into four main parts, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning, as shown in Fig. 1. Since there are many studies targeted at centralized unlearning, we use two sections to introduce related work in this scenario. We briefly classify the related unlearning methods into two main kinds, Exact unlearning (Cao and Yang, 2015; Bourtoule et al., 2021) and Approximate unlearning (Chien et al., 2022; Nguyen et al., 2020), where exact unlearning extends the idea of naive unlearning to design efficient retraining scheme with less computation cost. Approximate unlearning aims to unlearn a posterior that is approximate to the model retrained on the remaining dataset, only using the original learned model and the erased data. We then detailedly introduce the popular implementing techniques of these two unlearning kinds. In addition to these centralized unlearning methods, we introduce the distributed and irregular data unlearning scenarios, including federated unlearning (Wang et al., 2022; Liu et al., 2021), and graph unlearning (Chien et al., 2022; Chen et al., 2021a). After implementing unlearning, verifying the unlearning effect is equally important and draws lots of research attention. Therefore, we review the studies about unlearning verification subsequently. Moreover, we consider the privacy and security issues essential in machine unlearning, and we organize and review the related unlearning literature about privacy threats, defences, and unlearning applications in the survey.
There is no comprehensive survey about machine unlearning because it is a pretty new research domain. Only several reviews on the specific unlearning scenarios were published on Arxiv. An introduction to machine unlearning, about the exact and approximate unlearning problems and how they are defined and solved by recently proposed methods, see (Mercuri et al., 2022). About certifiable machine unlearning for linear models, the algorithm introduction and experiments analysis can be found in (Mahadevan and Mathioudakis, 2021). Nguyen et al. (Nguyen et al., 2022b) summarized the general unlearning framework and added the unlearning verification part to it. They paid most of their attention to introducing problem formulations and technical definitions. None of them noticed and reviewed the privacy and security issues in machine unlearning. Moreover, they have not comprehensively discussed the differences and connections between different unlearning scenarios.
In our treatment, we review not only centralized machine unlearning but also other scenarios, including distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning. We discuss these unlearning scenarios in the literature and compare the difference and connections from each other. For all four unlearning scenarios, we list open questions that, if solved, would increase our understanding of machine unlearning or enhance our ability to design a suitable unlearning mechanism for different applications.
The main contributions of the paper are as follows.
  • We systematically catalogue machine unlearning studies according to their scenarios, which can be briefly classified into, centralized unlearning, distributed and irregular data unlearning, unlearning verification, and privacy and security issues in machine unlearning.
  • For the centralized unlearning scenario, which draws the most unlearning research attention, we further divide related studies into exact and approximate unlearning and illustrate the connections, pros, and cons among these works.
  • We further explore the privacy threats that machine unlearning brings and discuss the applications of unlearning in defending against traditional security and privacy issues.
  • We discuss different challenges between various machine unlearning scenarios, from centralized unlearning to federated unlearning, graph unlearning, and privacy and security issues in unlearning. We list the related open questions of each scenario and present potential research directions to solve them.
The rest of the survey is organized as follows. The background of machine unlearning and the related technical tools are summarized in Section 2. After introducing the background, we will introduce the survey following the taxonomy order in Fig. 1. In Section 3, we present the primary categories of centralized unlearning: exact unlearning and approximate unlearning. In Section 4, techniques of centralized unlearning in different categories introduced in Section 3 are discussed in-depth. We collect and introduce the distributed and irregular data unlearning scenarios that we noticed, which are federated unlearning and graph unlearning, in Section 5. Section 6 introduces the unlearning evaluation and verification methods. The privacy and security issues in machine unlearning are divided into two sections. Section 7 introduces the privacy threats accompanied by machine unlearning. In Section 8, we discuss machine unlearning applications, which are mainly applied in dealing with security issues. In Section 9, we discuss the challenges of current unlearning methods in a general way. After that, we list the open gaps that new studies can target. At last, we summarise the survey in Section 10.
2. Background
2.1. Machine Unlearning
There are many urgent demands that motivate the hot machine unlearning research. The first is the privacy-preserving demand. Since the right to be forgotten has been legislated worldwide, machine unlearning is often formulated as a privacy-preserving problem where users can ask for their data erasure from the ML models. Besides privacy reasons, many other reasons promote the development of machine unlearning. For example, one common reason is usability. During the long time usage period of a recommendation application, users’ information will sometimes change, such as moving changes in their location. The application will produce unsuitable recommendations if it cannot completely delete the incorrect data from the model. Another reason is security. Deep learning models are vulnerable to adversarial attacks (Yuan et al., 2019) such as data poisoning (Tian et al., 2022). Hence, how to detect and remove adversarial poisoned data is essential to ensure models’ security. Once the attack type is detected, the model must erase these adversarial data’s influence using the unlearning mechanism.
Figure 2. Machine unlearning process.
To better understand how unlearning mechanisms work, we first introduce the unlearning problem and process following the machine unlearning framework demonstrated in Fig. 2. Let $\mathcal{Z}$ denote as a space of data items, then the particular (full) training dataset is $D\in\mathcal{Z}$. A learning process can be demonstrated as Step 0 in Fig. 2, i.e., training a model $M$ using an algorithm $\mathcal{A}$ on the training dataset $D$, denoted as $M=\mathcal{A}(D)$, where model $M$ is in a hypothesis space $\mathcal{H}$.
The unlearning process begins at an unlearning request when a user wants to remove his data $D_{e}$ from the trained model, Step 1 in Fig. 2. The requested unlearning data $D_{e}$ can be data samples, classes or graph nodes. Then, in Step 2, the server removes the contribution of $D_{e}$ using a machine unlearning algorithm $\mathcal{U}$. The unlearned model can be expressed as $M_{D\backslash D_{e}}=\mathcal{U}(D,D_{e},\mathcal{A}(D))$. In dealing with an unlearning request, we hope that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is the same as the retrained model $\mathcal{A}(D\backslash D_{e})$ (i.e., $\mathcal{U}(D,D_{e},\mathcal{A}(D))\simeq\mathcal{A}(D\backslash D_{e})$). Most unlearning studies ended at this step. However, we do need an effective evaluation metric to assess the unlearning effect as Step 3 in Fig.2. Therefore, we add the verification as the last step in the unlearning process, and we will introduce the related literature later.
2.2. Unlearning Evaluation Metrics
To compare the two models before and after unlearning, we need to define an evaluating metric $d(.)$ between $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ from the model parameter space $\mathcal{H}$. To this end, we briefly introduce several common evaluation distances metrics:
$L_{2}$-Norm. Wu et al. (Wu et al., 2020) proposed to utilise the Euclidean norm to evaluate the parameters of $\mathcal{A}(D\backslash D_{e})$ and that of unlearning $\mathcal{U}(D,D_{e},\mathcal{A}(D))$. Suppose $\theta$ are the model parameters learned by algorithm $\mathcal{A}(.)$. The $l_{2}$-norm or distance can be denoted as $||\theta_{\mathcal{A}(D\backslash D_{e})}-\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}||$, where $\theta_{\mathcal{A}(D\backslash D_{e})}$ is the model parameters of retraining from the scratch and $\theta_{\mathcal{U}(D,D_{e},\mathcal{A}(D))}$ is the modal parameters of unlearning algorithms $\mathcal{U}(.)$.
Kullback–Leibler divergence (KLD). In Bayes-based or Markov chain Monte Carlo-based unlearning methods, researchers used the Kullback–Leibler divergence or Jensen-Shannon divergence to optimize the approximate posterior since they can measure the distance of two probability distributions. Therefore, in recent unlearning studies, researchers also used the KLD between the retrained and unlearned models to estimate the unlearning effect (Nguyen et al., 2020). For discrete probability distritbutions of two model parameters $\mathcal{A}(D\backslash D_{e})$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ on the same probability space $\mathcal{H}$, the KLD from $\mathcal{A}(D\backslash D_{e})$ to $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ is denoted as $\text{KL}[\Pr(\mathcal{A}(D\backslash D_{e}))||\Pr(\mathcal{U}(D,D_{e},\mathcal{A}(D)))]$.
Parameter Leakage. Since membership inference attacks can infer if a sample is used in the training dataset of a model, recently, some works have leveraged the property to verify if the specific data is removed by unlearning mechanisms. Some studies (Hu et al., 2022) even proposed to backdoor the erased samples and then attack the unlearned model to check if the unlearned model can be attacked. If the unlearned model is backdoored, it proves that the unlearning algorithm cannot unlearn samples effectively. Conversely, if the backdoor trigger cannot attack the unlearned model, it proves that the unlearning algorithm is effective. Similar methods were also used in (Thudi et al., 2022; Ma et al., 2022) to evaluate the unlearning effect.
2.3. Majority of Tools Used in Unlearning
2.3.1. Differential Privacy (DP)
Differential privacy has become a popular standard for privacy protection in the statistic. In a DP model, a trusted analyzer collects users’ raw data and then executes a private mechanism to achieve differential privacy. The privacy requirement ensures the indistinguishability for any two outputs of neighbouring datasets, where neighbouring datasets mean the dataset only differs by replacing one user’s data, denoted as $X\backsimeq X^{\prime}$. A ($\epsilon,\delta$)-differential privacy algorithm $\mathcal{M}:\mathbb{X}^{n}\to\mathbb{Z}$ means that for every pair of neighboring datasets $X\backsimeq X^{\prime}\in\mathbb{X}^{n}$ and every subset $S\subset\mathbb{Z}$ has that $\mathcal{M}(X)\in S$ and $\mathcal{M}(X^{\prime})\in S$ are $\epsilon$-indistinguishable and $\delta$-approximate. The smaller the $\epsilon$ is, the more the privacy protection level increases. When the value of $\epsilon$ is 0, it implies that the outputted probability distribution of algorithm $\mathcal{M}$ cannot reflect any useful information about the datasets. A general DP mechanism based on adding Laplace noise was presented and theoretically analyzed in (Dwork et al., 2006).
2.3.2. Bayesian Variational Inference (VI)
In machine learning, the bayesian variational inference is used to approximate difficult-to-compute probability densities via optimization. In this part, we revisit the VI framework (Blei et al., 2017) that learns approximate posterior model parameters $\theta$ using Bayesian Theory (Bernardo and Smith, 2009). Given a prior belief $p(\theta)$ of an unknown model and a complete data trainset $D$, an approximate posterior belief $q(\theta|D)\sim p(\theta|D)$ can be optimized by minimizing the Kullback–Leibler divergence (KLD) (Joyce, 2011), $\text{KL}[q(\theta|D)||p(\theta|D)]$. The KLD is a type of statistical distance to measure how one probability distribution $q(\theta|D)$ differs from another probability distribution $p(\theta|D)$. However, it is intractable to compute the KLD exactly or minimize the KLD directly. Instead, the evidence lower bound (ELBO) (Blei et al., 2017) was proposed to be maximized, which is equivalent to minimize KLD between the two probability distributions. ELBO follows directly from $log(p(D))$ subtracting $\text{KL}[q(\theta|D)||p(\theta|D)]$, where the log marginal likelihood $log(p(D))$ is independent of $q(\theta|D)$. Since $\text{KL}[q(\theta|D)||p(D|\theta)]\geq 0$, the ELBO is a lower bound of $log(p(D))$.
In general training situations, ELBO is maximized using stochastic gradient ascent (SGA). The primary process is approximating the expectation $\mathbb{E}_{q(\theta|D)}[log(p(D|\theta))+log(p(\theta)/q(\theta|D))]$ with stochastic sampling in each iteration of SGA. We can use a simple distribution (e.g., the exponential family)to approximate computational ease posterior belief $q(\theta|D)$. When dealing with complex distribution whose density cannot be evaluated, we can maximize the ELBO with adversarial training by alternating between estimating the log-density ratio $log(p(\theta))/q(\theta|D)$ and maximizing the ELBO (Yu et al., 2019).
2.3.3. Privacy Leakage
Figure 3. Privacy leakage: a privacy reconstruction process
Privacy leakage occurs in both unlearning verification and privacy threats in unlearning two parts. In unlearning verification, researchers tried to use privacy leakage attacks to verify if the specific data is unlearned. In the privacy and security issues in unlearning, researchers tried to design effective inference attacks tailored to machine unlearning. The basic attack of privacy leakage in a machine learning setting is membership inference, which determines if a data sample was used in the ML model updating process or not. When a sample is fully known to the adversary, learning which model was trained based on it will leak information about the model. A generic membership inference process was introduced in (Shokri et al., 2017). Shokri et al. first trained the shadow models to simulate the target machine-learning model. Then, they observed and stored the different outputs of the shadow models based on different inputs, in or not, in the trainset. They used these stored outputs as samples to train the membership inference attack model.
Model inversion (Fredrikson et al., 2014), or privacy reconstruction (Salem et al., 2020) is another privacy threat in general machine learning. The goal of model inversion is to infer some missing attributes of input features based on the interaction with a trained machine learning model. Salem et al. (Salem et al., 2020) proposed a reconstruction attack target recovering specific data samples used in the model updating process by the different outputs of the model before updating and after updating. Latter, inferring the private information of updating data in conventional machine learning is transferred to infer the privacy of the erased samples in machine unlearning. In reconstruction attacks, the adversary first collects the different outputs using his probing data $D_{probe}$, including the original outputs $\hat{Y}_{\mathcal{M}}$ before unlearning, and the outputs $\hat{Y}_{\mathcal{M}_{-D_{e}}}$ after unlearning. Then he constructs the attack model based on the posterior difference $\delta=\hat{Y}_{\mathcal{M}_{-D_{e}}}-\hat{Y}_{\mathcal{M}}$. The attack model contains an encoder and decoder, which has a similar structure as VAEs (Kingma and Welling, 2014), and the main process is shown in Fig. 3.
3. Categories of Centralized Machine Unlearning
In this section, we classify existing centralized unlearning methods by their inherent mechanism and designed purposes. Detailed techniques will be introduced in the next section. From the former introduction, we know that naive retraining is the most effective manner to realize machine unlearning. However, it is inefficient because it needs to store the whole training dataset and retrain a new model from scratch. It consumes a huge amount of storage and computation costs, especially in deep learning scenarios. Therefore, researchers tried to design effective and efficient unlearning mechanisms. However, there were difficult challenges when we tried other unlearning methods. Three representative challenges are the stochasticity of training, the incrementality of training, and the catastrophe of unlearning. Exact unlearning was proposed to reduce the computation cost by splitting training sub-models based on pre-divided data sub-sets. It also divides the stochasticity and the incrementality into sub-models. It can unlearn an exact model by ensembling the consisting submodels, but it still needs to store all the split data subsets for fast retraining. Another approach is approximate unlearning. It can reduce both storage and computation costs because it unlearns only based on the erased data. However, it is difficult for approximate unlearning methods to control the accuracy degradation, due to the challenges in estimating stochasticity and incrementality during the training process. Most of them bounded their estimation to avoid dramatic accuracy reduction.
3.1. Exact Unlearning
Exact unlearning could also be called fast retraining, whose basic idea is derived from naive retraining from scratch. Following the background introduction of unlearning, we know the learning model $\mathcal{A}$ and unlearning model $\mathcal{U}$ based on the trainset $D$ and erased dataset $D_{e}\subseteq D$. If $\mathcal{U}(.)$ is implemented as naive retraining, the equality between $\mathcal{A}(D\backslash D_{e})\in\mathcal{H}$ and $\mathcal{U}(D,D_{e},\mathcal{A}(D))\in\mathcal{H}$ is absolutely guaranteed. However, naive retraining involves high computation and storage costs, especially for deep learning models and complex datasets (Thudi et al., 2022). Unlike naive retraining needs to retrain based on the whole remaining dataset, exact unlearning tries to retrain a sub-model based on the subset of remaining data to reduce computation cost. A general operation of exact unlearning is that they first divide the training dataset into several sub-sets. Then, they transform the learning process by ensembling the sub-models trained based on each sub-set as the final model. So that when unlearning request comes, they only need to retrain the trained sub-model based on the sub-set containing the erased data. They then ensemble the retrained sub-model and other sub-models as the unlearned model.
As we introduced above, we know that the primary goal of exact unlearning is to reduce the computation cost of retraining a new model by transforming the original learning algorithms into an ensembling form. It divides the stochasticity and incrementality into several sub-models to reduce their influence. However, to some extent, they sacrificed the storage cost because they needed to store the whole training dataset in a divided form. In (Cao and Yang, 2015), Cao and Yang transformed the traditional ML algorithms into a summation form. When an unlearning requirement comes, they only need to update several summations, ensuring the process runs faster than retraining from scratch. Bourtoule et al. proposed SISA (Bourtoule et al., 2021), which is short for Sharded, Isolated, Sliced, and Aggregated training. SISA divided the full data into shards and trained models separately in each shard. For unlearning, they simply need to retrain the shard that includes the erased data. Wu et al. (Wu et al., 2022a) proposed a framework that precisely models the influence of each training data point on the model concerning various performance criteria and removes the influence of the data that is required to delete. Golatkar et al. (Golatkar et al., 2021) proposed an unlearning method on deep networks, splitting the trained model into two parts. The core part based on the data will not delete, and the unlearning part on the erased data with parameters bound. These methods are efficient in computation, but they sacrifice the storage space to store the intermediate training parameters of different slices and the related training sub-sets.
Besides the huge storage cost, another big problem in exact unlearning is that it only suits the situation, in which the unlearning request only needs to remove a few samples with a low frequency. Suppose an unlearning request needs to remove many data samples (usually, they are not in the same previous divided sub-set). In that case, exact unlearning must retrains all these related sub-models or even all the sub-models in the worst situation. At the moment, exact unlearning is no longer computation efficient, and it still needs to store the whole training dataset or intermediate parameters.
3.2. Approximate Unlearning
Figure 4. Model changes when adding a new point or removing a point. (a) A normally trained classifying model classifies classes 1 and 2. (b) When a new point appears, the model is trained based on it, and the classifying line is pushed to classify it. (c) When we need to remove an influential point, we should recover the contribution of this point to the model. (d) When we remove a Non-influential point, the model may not need to change a lot.
Unlike exact unlearning could only reduce the retraining computation cost, approximate unlearning tries to directly unlearn based on the learned model and the erased data sample, which saves the computation and storage costs together. The goal of approximate unlearning is to unlearn a posterior that approximates the model trained on the remaining dataset, based on only the erased sub-dataset while not the remaining dataset (used by exact unlearning).
As we introduced before, the unlearning purpose is expecting that the unlearned model $\mathcal{U}(D,D_{e},\mathcal{A}(D))$ to equal to the retrained model $\mathcal{A}(D\backslash D_{e})$. Since exact unlearning is implemented by retraining from the remaining dataset or sub-sets, they can almost guarantee equality before and after unlearning. However, since approximate unlearning tries to directly delete the contribution of the erased data from the trained model, how to exactly estimate the contribution of the erased data (the contribution contains the stochasticity and the incrementality ) and remove it becomes the core problem.
The text description of the changes between two different distribution spaces before and after removing the specific data is not intuitive. Fig. 4 shows illustrated changes when adding a new point or removing a point in a classifying model. When an influential point appears, it usually pushes the line to move forward than the original classifying line to identify it, as shown in Fig. 4 (b). When this influential point is requested to remove, the unlearning mechanism must recover the model to the original one that has not been trained by this specific point, as shown in Fig. 4 (c). However, when only unlearning a non-influential point, which may have almost non-influence on the model, the unlearned model may not change compared to the original trained model in this situation, as shown in Fig. 4 (d).
Many methods were proposed to implement approximate unlearning efficiently and effectively. The popular solutions are certified-removal and Bayes-based mechanisms, which will be introduced in technical detail in Section 4. Here, we briefly summarize current studies about approximate unlearning. A certified-removal mechanism (Guo et al., 2020) was an approximate unlearning method similar to differential privacy. Guo et al. (Guo et al., 2020) defined the $\epsilon-$approximate unlearning, which makes sure that the model after and before unlearning must be $\epsilon$-indistinguishable as the definition in DP. The difference between $\epsilon-$approximate unlearning and $\epsilon$-DP is that the mechanism $\mathcal{A}$ on differential privacy is needed never to memorize the data in the first place, which is impossible in machine unlearning. The machine learning model does not learn anything from the training dataset if $\mathcal{A}$ is differentially private (Bourtoule et al., 2021). Sekhari et al. (Sekhari et al., 2021) introduced a similar solution to (Guo et al., 2020), which removes the contribution of the erased data from the gradients during the unlearning process. To mitigate the unlearning accuracy reduction, they bounded the model parameters by their differential unlearning definition. Neel et al. (Neel et al., 2021) proposed unlearning methods allowing arbitrary sequences of updates, including addition and deletion, and defined the corresponding weak and strong unlearning algorithms.
Different from indistinguishable unlearning, other methods used the Bayes theorem to unlearn an approximate posterior directly. In (Nguyen et al., 2020), Nguyen et al. unlearned an approximate posterior based on the erased data via the Bayesian inference. Similarly, Fu et al. (Fu et al., 2022) proposed a machine unlearning algorithm for Markov chain Monte Carlo (MCMC) and further defined a knowledge removal estimator the evaluate the knowledge erasing. Nguyen et al. (Nguyen et al., 2022a) also explained the effectiveness of the approximate machine unlearning algorithm from the perspective of Markov Chain Monte Carlo. However, since those methods are approximately unlearning the contribution of all the erasing data, including the inputs and labels, it is easy to decrease the model accuracy to some extent after unlearning.
4. Techniques of Centralized Unlearning
Table 1. Summary of popular unlearning techniques
Unlearning Literature Taxonomy Requests Type Techniques Realization Method Year
SISA (Bourtoule et al., 2021) Exact unlearning Samples Split learning Data and Model Partition 2021
GraphEraser (Chen et al., 2021a) Exact unlearning Graph nodes Split learning Data Partition 2021
Amnesiac unl. (Graves et al., 2021) Exact unlearning Samples Split learning Partially retraining 2021
RecEraser (Chen et al., 2022) Exact unlearning Samples Split learning Balanced Data Partition 2022
ARCANE (Yan et al., 2022) Exact unlearning Samples Split learning Parition by Class 2022
HedgeCut (Schelter et al., 2021) Exact unlearning Samples Split learning Tree ensemble learning 2021
DeltaGrad (Wu et al., 2020) Exact unlearning Samples Split learning L-BFGS (Byrd et al., 1994) 2020
L-CODEC (Mehta et al., 2022) Approximate unlearning Samples Certified Data Removal Markov Blanket selection 2022
PUMA (Wu et al., 2022a) Approximate unlearning Samples Certified Data Removal SME 2022
Certified Removal (Guo et al., 2020) Approximate unlearning Samples Certified Data Removal LP 2019
$(\epsilon,\delta)$-unl. (Neel et al., 2021) Approximate unlearning Samples Certified Data Removal Perturbed gradient descent 2021
(Ginart et al., 2019) Approximate unlearning Samples Certified Data Removal Quantized and DC-k-means 2019
Graph unl. (Chien et al., 2022) Approximate unlearning Graph nodes Certified Data Removal Certified removal 2022
EUBO, rKL (Nguyen et al., 2020) Approximate unlearning Samples Bayesian-based Unlearning VBI 2020
MCU (Nguyen et al., 2022a) Approximate unlearning Samples Bayesian-based Unlearning Monte cario-based 2022
BIF (Fu et al., 2022) Approximate unlearning Samples Bayesian-based Unlearning MCMC 2022
  • Alogrihtm abbreviations
  • VBI: Variational Bayesian Inference, SME: Store medial estimation, MCU: Monte Cario-based machine unlearning, FIM: Fisher Information Matrix, MCMC: Markov chain Monte Carlo, BIF: Bayesian inference forgetting, LP: Loss perturbation (Chaudhuri et al., 2011), TF-IDF: Term Frequency Inverse Document Frequency, EUBO: Evidence upper bound, rKL: reverse Kullback–Leibler
The naive retraining method is the most effective way to realize machine unlearning. However, it is expensive in both computation and storage. Exact unlearning extends the idea of retraining and tries to reduce the computation cost of unlearning. Approximate unlearning was proposed to find a way to reduce computation and storage consumption together. It is usually realized by unlearning based on the erased data only. In this section, we will introduce the popular techniques used in existing unlearning methods, in both exact unlearning and approximate unlearning. The dominant studies are summarized in Table 1, where the primary technique used in exact unlearning is split learning, and two primary techniques used in approximate unlearning are certified data removal and bayesian-based unlearning, respectively. The introduction order of these techniques will mainly follow the order of the unlearning techniques in Fig. 1.
4.1. Split Unlearning
Since most of the exact unlearning methods tried to split the full dataset into multi sub-sets and divide the model training process, we call this kind of unlearning technique split unlearning. The main process of split unlearning is illustrated in Fig. 5 (b). By contrast, we show the process of the naive retraining method in Fig. 5 (a). From Fig. 5 (a), we can see that when we unlearn the erased data by naive unlearning, we need two steps to realize it: first, remove the data from the original training dataset; second, retrain a new model based on the remaining dataset. Since it needs to store the whole training dataset and retrain from the remaining dataset, it often entails significant computational and stored overhead. Researchers saw the huge computation cost of naive retraining. They proposed many exact unlearning methods, and we summary as split unlearning because they were all based on data or model split techniques. As shown in Fig. 5 (b), the split unlearning technique can be summarized into four steps. Different from naive unlearning trained based on the whole remaining trainset, the first phase of split unlearning is dividing the original full trainset into multiple disjoint shards. All the constituent models are trained based on each split data slice. Then, in the second phase, when the unlearning request comes, they only need to remove the requested data from the split slice and retrain this slice’s constituent model in the third phase. In the last phase, the split unlearning aggregates the retrained and other constituent models together as a new unlearned model.
The first split unlearning is proposed by Cao and Yang (Cao and Yang, 2015). They split the original learning algorithms into a summation form. In a regular machine learning form, the model directly learns from the training dataset. However, in the summation form, they first train a small number of constituent models, which learn from several parts of the full trainset and then aggregate these intermediate models as the final learning model. So that when unlearning, they only need to retrain the constituent model that contains the information of erased data. It can efficiently speed up retraining time and reduce computation costs. In (Cao and Yang, 2015), the authors indicated that support vector machines, naive Bayes classifiers, k-means clustering, and many machine learning algorithms could be implemented in a summation form to reduce the retraining cost. The statistical query (SQ) learning (Kearns, 1998) guarantees the summation form. Although algorithms in Probably Approximately Correct (PAC) setting can be converted to the SQ learning setting, a huge amount of complex models, such as DNNs, cannot be converted to SQ learning efficiently.
Figure 5. (a) Naive unlearning. Only two steps: remove the erased data from the whole dataset and retrain the model based on the remaining dataset. (b) Split unlearning. It contains four steps: 1. split the original dataset into $n$ shards, 2. remove the erased data from the corresponding shard, 3. retrain the sub-model of this shard, 4. ensemble all sub-models as the final model.
Then, Bourtoule et al. (Bourtoule et al., 2021) and Yan et al. (Yan et al., 2022) proposed advantaged methods unlearn samples suitable on deep neural networks. The primary idea of (Bourtoule et al., 2021; Yan et al., 2022) is also similar to the process shown in Fig. 5 (b). In (Bourtoule et al., 2021), Bourtoule et al. named their unlearning method the SISA training approach. SISA was implemented on deep neural networks, including two Conv. layers followed by two FC layers, Wide ResNet-1-1 and ResNet-50 models on multiple datasets, including MNIST, Purchase, SVHN, CIFAR100, Imagenet, and Mini-Imagenet. Unlike the original split unlearning dividing the dataset and transforming learning algorithms to summation form, Yan et al. proposed ARCANE (Yan et al., 2022), which transforms conventional ML into ensembling multiple one-class classification tasks. When many unlearning requests come, it can reduce retraining costs, which is not considered by previous work.
Chen et al. extended exact unlearning methods to recommendation tasks and proposed RecEraser (Chen et al., 2022), which has similar architecture as split unlearning in Fig. 5 (b). RecEraser is tailored to recommendation systems, which can efficiently implement unlearning. Specifically, they designed three data partition schemes to split recommendation data into balanced pieces and propose an adaptive aggregation method based on an attention mechanism. They conducted the experimental evaluation on three real-world datasets, YELP2018, Movielens-1m, and Movielens-10m. These three datasets are widely used to assess the effectiveness and efficiency of recommendation models.
Schelter et al. proposed HedgeCut (Schelter et al., 2021), which implemented machine unlearning on tree-based ML models in a split unlearning similar form. Tree-based learning algorithms are trained by recursively partitioning the training data, locally optimizing a criterion such as Gini gain (Wehenkel and Pavella, 1991). HedgeCut focuses on implementing fast retraining for these methods. Furthermore, they evaluated their method on five publicly available datasets on both accuracy and running time.
Another method that is similar to split unlearning is Amnesiac Unlearning (Graves et al., 2021). The intuitive idea of Amnesiac Unlearning is to store the parameters of training batches and then subtract them when unlearning requests appear. In particular, it first trains the learning model by adding the total gradients $\sum_{e=1}^{E}\sum_{b=1}^{B}\nabla_{\theta_{e,b}}$ to the initial model parameters $\theta_{initial}$, where $E$ is the training epochs, and $B$ is the data batches. During the training process, they kept a list of $SB$, which records the batches containing the private data. The list could be in the form of an index of batches for each training example, an index of batches for each training category or any other information expected. When the unlearning request comes, a model using Amnesiac unlearning needs only to remove the updates from each batch $sb\in SB$ from the learned model $\theta_{M}$. As Graves et al. mentioned, using Amnesiac unlearning effectively and efficiently removes traces of the erased data that could be inferred via state-of-the-art privacy inference attacks and does not degrade the accuracy of the model in any other way.
4.2. Certified Data Removal
Certified data removal unlearning methods usually define their unlearning algorithms as $\epsilon$-indistinguishable unlearning, which is similar to the differential privacy definition. Most of them use the Hessian matrix (Bishop, 1992) to evaluate the contribution of erased data samples for unlearning subtraction. After estimating the contribution of the erased data samples, they unlearn by subtracting these contributions with an updating bound from the unlearning model.
In (Guo et al., 2020), Guo et al. proposed a certified data removal mechanism, which assumes removing the last training sample, $(x_{n},y_{n})$. Specifically, they defined a removal mechanism that approximately minimizes $\mathcal{L}(\theta;D^{\prime})$ with $D^{\prime}=D\backslash(x_{n},y_{n})$. The loss gradient at sample $(x_{n},y_{n})$ can be denoted as $\Delta=\lambda\theta^{*}+\nabla l((\theta^{*})^{T}x_{n},y_{n})$ and the Hessian of $L(\cdot;D^{\prime})$ at $\theta^{*}$ by $H_{\theta^{*}}=\nabla^{2}L(\theta^{*};D^{\prime})$. Then they applied a one-step Newton update to the gradient impact of the removed point $(x_{n},y_{n})$ on the model $\theta$. Under their observation, they found that directly removing the Hessian contribution from the direction of the gradient will leak the privacy of the training data that was expected to be removed. They used the loss perturbation technique (Chaudhuri et al., 2011) to hide this information. It used a random linear term to perturb the empirical risk and ensures that the outputs of their method $\mathcal{U}(D,D_{e},\mathcal{A})$ is $\epsilon$-indistinguishable between the retrained model $\mathcal{A}(D\backslash D_{e})$. In(Mehta et al., 2022) and (Thudi et al., 2022), they designed unlearning algorithms following the certified data removal definition in (Guo et al., 2020). Mehta et al. unlearned via their proposed efficient Hessians, L-FOCI (Mehta et al., 2022). Thudi et al. (Thudi et al., 2022) used membership inference as a verification error to adjust the unlearning process on stochastic gradient descent (SGD) optimization.
Another similar unlearning method is PUMA. In (Wu et al., 2022a), Wu et al. proposed a new data removal method through gradient re-weighting called PUMA, which also used the Hessian Vector Product (HVP) term. They first estimated and recorded individual contributions of $(x_{i},y_{i})$, where the estimation is limited to less than a dot product between the pre-cached HVP term and individual gradient. When the unlearning request comes, they subtract the estimation of the erased samples to update the model parameters.
Ginart et al. (Ginart et al., 2019) extended certified data removal to k-means clustering algorithms. They formulated the unlearning problem of efficiently removing personal data information from trained clustering models and proposed two different efficient deletions for k-means clustering, quantized k-mean and divide-and-conquer k-means. In their work, both algorithms have theoretical guarantees and strong empirical results.
DeltaGrad (Wu et al., 2020), an algorithm for fast retraining of SGD-based models, was proposed by Wu et al. to small changes of data, based on the idea of ”differentiating the optimization path” concerning the data and Quasi-Newton methods. They theoretically proved that their algorithm could approximate the right optimization path rapidly for the strongly convex objective. DeltaGrad starts with a ”burn-in” period of first iterations, where it computes the full gradients exactly. After that, it only calculates the complete gradients for every first iteration. For other iterating rounds, it operates the L-BGFS algorithm (Byrd et al., 1994) to compute Quasi-Hessians approximating the true Hessians, keeping a set of updates at some prior iterations.
Sekhari et al. (Sekhari et al., 2021) further given a strict separation between $\epsilon$-indistinguishable unlearning and differential privacy. Different from (Guo et al., 2020), in order to utilise tools of differential privacy (DP) for ML, the most straightforward manner is to ignore the special set of erasure demands $D_{e}$ and create an unlearning mechanism $\mathcal{U}$ that only relies on the learned algorithm $\mathcal{A}(D)$. In particular, the unlearning algorithm is of the form $\mathcal{U}(D_{e},\mathcal{A}(D))=\mathcal{U}(\mathcal{A}(D))$ and makes sure the true unlearned model $\mathcal{U}(\mathcal{A}(D))$ is $\epsilon$-indistinguishable to $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Notice the difference between (Guo et al., 2020) and (Sekhari et al., 2021). In the definition of (Guo et al., 2020), their $\epsilon$-indistinguishable unlearning is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{A}(D\backslash D_{e})$, but here is between $\mathcal{U}(\mathcal{A}(D))$ and $\mathcal{U}(\mathcal{A}(D\backslash D_{e}))$. Such a pair of algorithms in (Sekhari et al., 2021) would be differential private for $D$, where the neighbouring datasets mean that for two datasets with an edit distance of $m$ samples. The guarantee of DP unlearning is stronger than the model distribution undistinguishable unlearning guarantee in (Guo et al., 2020), and thus it suffices to satisfy it. Based on the definition of (Sekhari et al., 2021), they pointed out that any DP algorithm automatically unlearns any $m$ data samples if it is private for datasets with edit distance $m$. Therefore, they derivate the bound on deletion capacity from the standard performance guarantees for DP learning. Furthermore, they determine that the existing unlearning algorithms can delete up to $\frac{n}{d^{1/4}}$ samples meanwhile still maintaining the performance guarantee w.r.t. the test loss, where $n$ is the size of the original trainset and $d$ is the dimension of trainset inputs.
The above methods are all answering the problem of unlearning. Neel et al. (Neel et al., 2021) extended the unlearning definition to the update, which includes both ’add’, $D\cup\{z\}$, and ’delete’ as $D\backslash\{z\}$, where $z\in\mathcal{Z}$ is a data point. They follow the definition in (Guo et al., 2020) and define the similar $(\epsilon,\delta)$-indistinguishability. Moreover, they extend $(\epsilon,\delta)$-indistinguishability for both ’add’ and ’delete’ updates, which can also be denoted as $(\epsilon,\delta)$-publishing.
4.3. Bayesian-based Unlearning
Different from certified data removal that unlearns samples by subtracting corresponding Hessian matrix estimation from trained models, Bayesian-based unlearning tries to optimize an approximate posterior of the remaining dataset only using the trained model (the posterior of the original full dataset) and the erased data. The exact Bayesian unlearning posterior can be derivated from the Bayesian rule as $p(\theta|D_{r})=p(\theta|D)\ p(D_{e}|D_{r})/p(D_{e}|\theta)$, where $\theta$ is the posterior (i.e., model parameters). $D$ is the original full trainset, and $D_{r}$ is the remaining dataset which removes the erased data $D_{e}$, denoted as $D=D_{e}\cup D_{r}$ and $D_{e}\cap D_{r}=\emptyset$. If the model parameters $\theta$ are discrete-valued, $p(\theta|D_{r})$ can be directly obtained from the Bayesian rule. The use of a conjugate prior furthermore makes unlearning relatively simple.
However, in practice, it is not easy to achieve the exact posterior, not to mention the unlearning posterior. In (Nguyen et al., 2020), Nguyen et al. tried doing likewise at the beginning. They defined the loss function using the KLD between the approximate predictive distribution $q_{u}(y|D_{r})$ and the exact predictive distribution $p(y|D_{r})$ where the observation $y$ (i.e., drawn from a model with parameters $\theta$) is conditionally independent of $D_{r}$ given $\theta$. They bounded this loss function by the KLD between posterior beliefs $q_{u}(\theta|D_{r})$ and $p(\theta|D_{r})$ and further proposed evidence upper bound (EUBO) as the loss function to unlearn the approximate unlearning posterior. To avoid the overestimation of using KL divergence to optimize the posterior, they introduced an adjusted likelihood to control the unlearning extent.
In (Nguyen et al., 2022a), the authors also studied the problem of ” unlearning” particular erased subset samples from a trained model with better efficiency than retraining a new model from scratch. Toward this purpose, Nguyen et al. proposed a Markov chain Monte Cario-based machine unlearning (MCU) method deriving from the Bayesian rule. They experimentally proved that MCU could unlearn the erased subsets of the whole training dataset from a prepared model effectively and efficiently.
Fu et al. (Fu et al., 2022) converted the Markov chain Monte Carlo unlearning problem into an explicit optimization problem. Then they proposed $\epsilon-$Knowledge removal, which was a little similar to certified removal methods, but they defined that the KLD between the unlearned and the retrained model must be less than $\epsilon$. To quantify the explicit $\epsilon-$knowledge removal, they proposed a knowledge removal estimator to evaluate the difference between the original and unlearned distributions. As they indicated, though their algorithm cannot wholly remove the learned knowledge from the already trained distribution, their method can still help the unlearned model approach its local minimum.
5. Distributed and Irregular Data Unlearning
We have introduced popular centralized machine unlearning in the former two sections. Most of them are requested to delete the regularly structured data. Their main purpose is to reduce the computation and storage costs and the accuracy degradation caused by unlearning. Recently, we noticed that some researchers paid attention to realizing unlearning in other scenarios, such as in distributed and irregular data learning scenarios. In this section, we will introduce the literature about federated unlearning and graph unlearning as two representative directions.
5.1. Federated Unlearning
Federated learning (McMahan et al., 2016, 2017) is first proposed to protect the participating users’ privacy during a machine learning training process in distributed situations. All participants will only upload their locally trained model parameters instead of their sensitive local data to the server for the model training. Therefore, in a federated learning scenario, the limited access to the dataset will become a unique challenge to implement unlearning.
Since the local data cannot be uploaded to the FL server side, most federated unlearning methods tried to unlearn a certain client’s influence from the trained model by storing and estimating the contribution of uploaded parameters. In this situation, they can implement federated unlearning without interacting with the client. The two representative methods are (Liu et al., 2021; Wu et al., 2022b). Liu et al. (Liu et al., 2021) proposed FedEraser to sanitize the contribution of a federated client on the global federated learning model. In particular, during the FL global training process, the FL-Server maintains the updates of the clients at each routine iteration and the index of the related round to calibrate the retrained updates. Based on these operations, they reconstructed the unlearned global model instead of retraining a new model from scratch. However, FedEraser can only unlearn one client’s data, which means it must unlearn all the contributions of this specific client’s data. It is unsuitable for a client who wants to unlearn a small piece of his data. Wu et al. (Wu et al., 2022b) tried to unlearn a client’s influence from the federated global model by removing the historical updates from the global model. They implemented federated unlearning by using knowledge distillation to restore the contribution of clients’ models, which does not need to rely on clients’ participation and any data restriction.
Wang et al. (Wang et al., 2022) explored the problem of how to selectively unlearn a category from a trained CNN classification model in FL. The biggest challenge is that the local data used for FL global model’s training cannot be accessed globally. Therefore, they explored the inner influence of each channel in deep and observed that various channels have distinct impacts on different categories. They proposed a method that does not require accessing the data used for training and retraining from scratch to cleanly scrub the information of particular categories from the global FL model. A method called Term Frequency Inverse Document Frequency (TF-IDF) was introduced to quantize the class discrimination of channels. Then, they pruned the channels with high TF-IDF scores on the erased classifications to unlearn them. Evaluations of CIFAR10 and CIFAR100 demonstrate the unlearning effect of their method.
Different from unlearning a whole client’s influence and unlearning a class, Liu et al. (Liu et al., 2022a) investigated how to unlearn data samples in FL. They first defined a federated unlearning problem and proposed a fast retraining method to remove data points from the federated learning global model. Then, they proposed an efficient federated unlearning method following the Quasi-Newton methods (Xu et al., 2020) and the first-order Taylor approximate technique. They used the practical Fisher Information Matrix (FIM) (Ly et al., 2017) to model the Hessian matrix at a low cost. When unlearning, they subtract the estimation matrix from gradients with a bound.
5.2. Graph Unlearning
We introduce graph unlearning as a representative kind of irregular data unlearning. In (Chien et al., 2022; Chen et al., 2021a; Cong and Mahdavi, 2022), researchers extend regular data machine unlearning to a graph data scenario. Graph structure data are more complex than standard structured data because graph data include not only the feature information of nodes but also the connectivity information of different nodes, shown in Fig. 6. Therefore, Chien et al. (Chien et al., 2022) proposed node feature unlearning, edge unlearning, and both node and edge unlearning for simple graph convolutions (SGC). Besides the different information unlearned in a graph learning problem, they found another challenge associated with feature mixing during propagation, which needs to be addressed to establish provable performance guarantees. They gave the theoretical analysis for certified unlearning of GNNs by illustrating the underlying investigation on their generalized PageRank (GPR) extensions and the example of SGC.
Figure 6. (a) Regular data unlearning, which only unlearns the data sample. (b) Graph unlearning. Since the graph data contains both node features and connective edge information, we may need to unlearn only edges information or both edges and nodes in graph unlearning.
Chen et al. (Chen et al., 2021a) found that applying SISA (Bourtoule et al., 2021) unlearning methods to graph data learning will severely harm the graph-structured information, resulting in model utility degradation. Therefore, they proposed GraphEraser to implement unlearning tailored to graph data. Similar to SISA, they first cut off some connecting edges to split the total graph into some sub-graphs. Then they trained the constituent graph models on these sub-graphs and ensemble them for the final prediction task. To realize graph unlearning efficiently, they proposed two graph partition algorithms and corresponding aggregation methods based on them.
In (Cong and Mahdavi, 2022), Cong and Mahdavi filled in the gap between regularly structured data unlearning and graph data unlearning by studying the unlearning problem on the linear-GNN. To unlearn the knowledge of a specific node, they propose a projection-based unlearning approach PROJECTOR that projects the weight parameters of the pre-trained model onto a subspace that is irrelevant to the deleted node features. PROJECTOR could overcome the challenges caused by node dependency and is guaranteed to unlearn the deleted node features from the pre-trained model.
6. Unlearning Evaluation and Verification
Table 2. Evaluation and Verification Metrics, Datasets
Evaluation Metrics Formula/Description Usage Literature Popular Datasets
Accuracy Accuracy of the unlearned model on erased datasets and remaining datasets To evaluate the predictive accuracy of the unlearned model (Bourtoule et al., 2021; Guo et al., 2020), … MNIST (LeCun et al., 1998) CIFAR SVHN LSUN Photo (Shchur et al., 2018) ImageNet (Deng et al., 2009) Adult Credit info PID (Asuncion and Newman, 2007) PWD (Liu et al., 2010) AG News (Zhang et al., 2015) GTSRB (Stallkamp et al., 2011) Covtype (Blackard and Dean, 1999) HIGGS (Baldi et al., 2014) RCV1 (Glewis et al., 2004) Market-1501 (Zheng et al., 2015) YELP2018 Movielens-1m Movielens-10m Cora (Sen et al., 2008) Citseer (Yang et al., 2016) Pubmed (Fey and Lenssen, 2019) ogbn-arxiv (Hu et al., 2020) Computers (McAuley et al., 2015) CS (Shchur et al., 2018) Physics (Shchur et al., 2018)
Running time The training time of unlearning process To evaluate the unlearning efficiency (Cao and Yang, 2015; Guo et al., 2020), …
L2-norm The parameters differences between the retrained and unlearned models $||\theta_{1}-\theta_{2}||$ To evaluate the indistinguishability between two models (Wu et al., 2020)
KL-Divergence The KLD between the distribution of the unlearned and retrained model: $\text{KL}(A(D_{r})||U(D_{e},A(D)))$ To evaluate the indistinguishability between model parameters (Nguyen et al., 2020)
JS-Divergence The JS divergence between the predictions of retrained and unlearned model: $JS(A,U)=0.5\cdot KL(A||Q)+0.5\cdot KL(U||Q)$ To evaluate the indistinguishability between model outputs (Chundawat et al., 2022)
Membership inference Recall (##\##detected items /#absent#/\#/ #erased items) To verify if the erased data is unlearned by the model (Graves et al., 2021)
Forgetting rate (FR) $FR=\frac{AF-BF}{BT}$, where AF, BF and BT are defined below To measure the rate of samples that are changed from member to non -member after unlearning (Ma et al., 2022)
[…]
EMA Ensembled Membership Auditing: Ensemble multiple membership metrics and utilizes Kolmogorov-Smirno (KS) statistical tools to obtain a final auditing score To verify if a unlearned model mem -orizes a query dataset (Liu and Tsaftaris, 2020)
MIB Membership Inference via Backdooring: Achieve membership inference for the backdoored data by querying a certain number of black-box queries to the model To verify if the model unlearns the backdoored data (Hu et al., 2022)
Recent studies paid a huge amount of attention to unlearning problem-solving; however, verifying the unlearning effect is also an important problem in unlearning. In Section 2, we have introduced the L2-norm (Wu et al., 2020), KLD (Nguyen et al., 2020), and privacy leakage as the unlearning verification metrics. The common metrics also include the accuracy to assess the predictive accuracy of the unlearned model and running time to evaluate the unlearning efficiency. We have listed recent evaluation and verification metrics in Table 6, where the popular experimental datasets are furthermore presented. Since accuracy, running time, L2-norm, and KLD are fixed evaluation metrics, we do not discuss them in detail here. In this section, we pay attention to the new evaluation metrics that recent studies tailored to unlearning. Most of them are based on privacy leakage.
In (Liu and Tsaftaris, 2020), Liu and Tsaftaris indicated that the data auditing problem in machine unlearning is becoming popular. They used a calibration dataset and the Kolmogorov-Smirno (KS) (Demidenko, 2004) distance to confirm if an erased dataset is removed. Their challenge was that the method would fail when the calibration dataset was out of high quality or when the query samples were similar to the training dataset. An Ensembled Membership Auditing (EMA) (Huang et al., 2021a) method was proposed to overcome these limitations when auditing data erasure from a trained model. EMA is a 2-step data removal auditing method that ensembles several metrics and statistical tools. In the first step of EMA, the auditor uses various metrics to infer if the model remembers the samples of the query dataset. In the second step, EMA aggregates a final auditing score by ensembling many membership metrics and statistical tools. They verified whether a trained model forgets query data by the final auditing score. They conducted the experiments using standard datasets, including MNIST, SVHN, and the Chest X-ray dataset.
Inspired by backdoor attacks in ML, Hu et al. (Hu et al., 2022) proposed Membership Inference via Backdooring (MIB). MIB leverage the property of backdoor attacks that backdoor triggers will misadvise the trained model to predict the backdoored sample to other wrong class. The main idea of MIB is that the user proactively adds the trigger to her data when publishing them online so that she can implement the backdoor attacks to determine if a model has been trained based on her data. MIB evaluate the membership inference for the triggered data by calculating the results of a certain number of black-box queries to the targetted model. Then, they observed the results between the targeted model and clean models to infer if the model was trained using the backdoored data.
A similar membership-inference-based method was proposed in (Ma et al., 2022). Ma et al. verified the effectiveness of an unlearning method by their proposed uniform metric called forgetting rate (FR). Their methods were inspired by membership inference attacks and defined the evaluation metrics based on the observation of membership inference. Suppose the dataset $D_{e}$ is the erased dataset; the FR of an unlearning method is denoted as $FR=\frac{AF-BF}{BT}$. In their FR definition, $BF$ and $AF$ mean the number of samples in $D_{e}$ that are predicted to be false by a membership inference attack before and after unlearning. $BT$ is the number of samples in $D_{e}$ predicted to be true by a membership inference attack before machine unlearning. According to this definition, $FR$ presents an instinctive evaluation of how many data points are changed from member to non-member after unlearning. If an unlearning method achieves that $AF>BF$ on the condition that $BT>0$, it means this method is effective. By contrast, the unlearning will be meaningless. They also proposed an unlearning method called Forsaken and evaluated Forsaken’s performance on eight standard benchmark datasets. The main idea of Forsaken is to continuously generate masked gradients, apply them to the model’s neurons, and model them to forget the contribution of the specific samples. To implement this process, they introduced a mask gradient generator. The proposed mask gradient generator could adjust the volume of the mask gradients to avoid unlearning catastrophes based on the observation of the updated model.
Sommer et al. (Sommer et al., 2022) introduced Athena, which leverages the property of backdoors, to verify the effectiveness of unlearning. Athena effectively certifies whether the data is deleted from an unlearning method with high confidence. Thus, it provides a basis for quantitatively inferring unlearning. In their backdoor-based verification scheme, they first backdoor users’ data and then test the backdoor success probability to infer if the data is deleted.
7. Privacy Threats on Machine Unlearning
Although machine unlearning was first proposed to protect users’ privacy, many researchers have noticed that unlearning brings new privacy threats simultaneously. Similar to machine learning scenarios, there are two main kinds of privacy threats in unlearning. They are membership inference attacks and private information reconstruction attacks. In this section, we introduce the two prevalent unlearning privacy threats detailedly.
7.1. Membership Inference Attacks in Unlearning
Chen et al. (Chen et al., 2021b) first pointed out that when a model is unlearned, the difference in the outputs from the model before and after unlearning leaks the privacy of the erased data. Then they proposed the corresponding membership inference attack pipeline in unlearning, which includes three phases: posteriors generation, feature construction, and membership inference.
  1. (1)
    Posteriors Generation. Suppose that the attacker has the ability to access two versions of the ML model, the model $\theta_{\mathcal{A}}$ before unlearning and the unlearned model $\theta_{\mathcal{U}}$. Assume a target sample $e$, the attacker queries $\theta_{\mathcal{A}}$ and $\theta_{\mathcal{U}}$, and obtains the corresponding posteriors, $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, which also called as confidence values in (Shokri et al., 2017).
  2. (2)
    Feature Construction. After achieving the two posteriors $p(\theta_{\mathcal{A}})$ and $p(\theta_{\mathcal{U}})$, the attacker sums them to make the inference feature vector $F$. Common methods exist to construct the feature vector shown in (Chen et al., 2021b).
  3. (3)
    Inference. After the attacker finishes training the attack model based on the created features $F$, he inputs the collected feature to the inference model to predict if the specific sample $e$ is in the erased dataset of unlearning models.
In (Chen et al., 2021b), they assumed that the attacker has admission to two versions of ML models before and after unlearning, but it is sometimes impractical, especially in black-box learning scenarios. Lu et al. (Lu et al., 2022) further proposed a label-only membership inference attack method to infer if a data point is unlearned by the model, eliminating the dependence on accessing posteriors. Their basic idea is that the same noise injection on candidate data points will show different results for the sample in or not in the training dataset. Thus, they made the adversary continuously query the original and unlearned models and add noise to change their outputing labels. By observing the disturbance amplitude, they can determine whether an item is deleted.
Golatkar et al. (Golatkar et al., 2020b) derived an upper bound to confirm the maximizing knowledge that can be extracted from a black-box model. They queried the model with an image and obtained the related output. They then used the entropy of the result probabilities to construct an effective black-box membership inference (Shokri and Shmatikov, 2015) attack in machine unlearning.
7.2. Privacy Reconstruction Attacks in Unlearning
Privacy reconstruction is another popular attacking kind in machine unlearning. In an unlearning scenario, Gao et al. (Gao et al., 2022) proposed the deleted reconstruction attacks to recover the removed data from the outputs of the original and unlearning models. In their work, they formalized erasure inference and erasure reconstruction attacks, in which the attacker aims to either infer which record is deleted or reconstruct the erased sample. In particular, for the deletion inference, they formalize the goal of an erasure inference to distinguish a data example $e$ was in or not in the erased dataset, $e\in D_{e}$ or $e^{\prime}\notin D_{e}$. For the deletion reconstruction, they focused on reconstructing the erased example $e$. In all their reconstruction attacks, the attacker is not given any particular examples, and its goal is to extract knowledge about the features of the erased instance. Specifically, the deletion reconstructions include the deleted instance reconstruction and deleted label reconstruction. As named, the deleted instance reconstruction is to extract all of the features of the erased instance, and the erased label reconstruction is to infer the label of the erased point in the classification problem.
Zanella-Béguelin et al. (Zanella-Béguelin et al., 2020) indicated that the releasing snapshot of overlapped language models would leak the privacy of the training dataset. They verified that the model updates pose a significant risk to the private information added to or removed from the training dataset by a huge amount of experimental results. Zanella-Béguelin et al. found five phenomenons. First, an attacker can extract particular sentences used or not in the training dataset by comparing two models. Second, analysing more model snapshots shows more information about the updated data than considering fewer model snapshots. Third, adding or deleting other non-private data during model updates can not mitigate privacy leakage. Fourth, differential privacy can reduce privacy leakage risks, but it also decreases the accuracy of trained models. Fifth, to mitigate the privacy leakage risks while keeping the model utility, the server can limit the model parameters access or only output a subset of the results.
Many studies furthermore utilised these privacy threats to evaluate the unlearning effect. Huang et al. (Huang et al., 2021a) proposed Ensembled Membership Auditing (EMA) for auditing data erasure. They use the membership inference to evaluate the removing effect of unlearning. Graves et al. (Graves et al., 2021) indicated that if an attacker can infer the sensitive information that was wanted to be erased, then it means that the server has not guarded the user’s rights to be forgotten. Baumhauer et al. (Baumhauer et al., 2022) developed linear filtration to sanitise classification models with logits prediction after class-wide deletion requests. They verified their methods by testing how well it defends against privacy attacks.
Both (Chen et al., 2021b) and (Zanella-Béguelin et al., 2020) pointed out that differential privacy guarantees that a model does not reveal too much information about any particular training data point. A differentially private model can furthermore guarantee the group’s privacy by bounding the impacts of a group of training samples. If using DP to protect the privacy of a group of $|D\backslash D_{e}|$training samples against snapshot attacks on $\theta_{D},\theta_{D\backslash D_{e}}$, it means that $\theta_{D\backslash D_{e}}$ cannot be more useful than $\theta_{D}$.
8. Machine Unlearning Applications
Table 3. Machine Unlearning Application
Literature Application Scenarios Algorithms Realization Methods Evaluation Metric Year
(Huang et al., 2021b) Data unlearnable EMP Add noise ACC 2021
(Liu et al., 2022b) Mitigate backdoor BAERASER Gradient ascent method ASR and ACC 2022
(Wang et al., 2019) Mitigate backdoor - Median Absolute Deviation L1-Norm, ##\##FP, ACC, ASR 2019
(Du et al., 2019) Anomaly detection Unlearning Maintain memory set ##\##FP and ##\##FN 2019
(Cao et al., 2018) Repair pollution KARMA Cluster and unlearn DA 2018
  • ASR:Attack Success Rate, ACC: Accuracy
  • DA:Detection Accuracy, EMP: Error-minimizing Perturbations
  • ##\##FP: number of false positive, ##\##FN: number of false negative
Besides the inherent demands in machine unlearning that draw much research attention, many researchers also find that machine unlearning can be applied in many other scenarios to solve related problems. We list the recent unlearning applications in Table 3. The most popular application of machine unlearning is to mitigate the anomaly, including the backdoor triggers and pollution, from a trained model.
Wang et al. (Wang et al., 2019) applied machine unlearning in detecting and mitigating backdoor attacks in DNNs. They designed two approaches to eliminate backdoor-related neurons/weights from the infected model and repair the infected model to be robust against malicious images. Then, Liu et al. (Liu et al., 2022b) proposed that a backdoor model learned the poisoned decision boundary. Data points with triggers are all classified into the target class. They reversed the backdoor injection process to defend against it in machine unlearning, which is simple but effective. Their method contains two primary steps: first, they use a max-entropy staircase approximator to complete trigger reconstruction; second, they remove the added backdoor triggers using unlearning. They named these two key steps of BAERASER as trigger pattern recovery and trigger pattern unlearning. Via a dynamic penalty mechanism, they mitigated the sharp accuracy degradation of gradient-ascent-based machine unlearning methods.
Repairing pollution is another successful unlearning application. Cao et al. (Cao et al., 2018) proposed KARMA to search different subsets of training datasets and return the subset that with the highest misclassifications. KARMA includes three parts. First, KARMA searches for potential causality that leads to the wrong classification of a ML model. It clusters the misclassified data points into different parts and extracts the centres of clusters. Then KARMA prioritizes the search for similar data samples in the training datasets based on these extracted centres. Second, KARMA grows the causality found in the first stage by finding more training samples and forming a cluster. Third, KARMA determines if a causality cluster is polluted and calculates how many samples the cluster contains.
Machine unlearning was also applied to lifelong anomaly detection problems. Du et al. (Du et al., 2019) proposed new approaches to solve the challenges of exploding loss and sharp accuracy degradation caused by unlearning. In their presentation, they explain their techniques, emphasizing handling false negatives. When meeting such case $x_{t}$, their goal is, therefore, to modify the model to reduce the probability of $p(x_{t})$, or to increase the loss $\mathcal{L}(x_{t})$.
Huang et al. (Huang et al., 2021b) presented a method that can make training examples unlearnable by injecting a type of error-minimizing noise. This kind of noise is intentionally generated to reduce the error of the training examples close to zero, which can mislead the model into considering there is ”nothing” to learn from these examples. They first tried the error-maximizing (Goodfellow et al., 2014) noise but found that this noise cannot prevent DNN learning when applied in a sample-wise manner to the training data points. Therefore, they then begin to study the opposite direction of error-maximizing noise. In particular, they proposed the error-minimizing noise to stop the model from being punished by the loss function during traditional ML model training. Therefore, it can mislead the ML model to consider that there is ”nothing” to learn.
9. Lessons Learnt and Discussions
In the research domain of machine unlearning, researchers mainly face three difficult challenges, the stochasticity of training, incrementality of training, and catastrophe of unlearning, when they try to design efficient and effective unlearning mechanisms replacing naive retraining. They tried many mechanisms to mitigate the influence of these challenges. For example, in exact unlearning, they designed split algorithms that divide the final model into several sub-models and avoid the stochasticity and incrementality of one sub-model to influence other sub-models. In approximate unlearning, they bounded the removed estimation to aviod the accuracy degradation. However, it is still not easy to analyze them clearly and solve them totally. Moreover, as the research in machine unlearning in-depth, researchers extended it to new situations, such as federated learning and graph learning, and they met new challenges. In this section, we will discuss the differences between various unlearning scenarios and their corresponding challenges. Based on these challenges, we list some potential research directions on which we can focus.
In the common centralized unlearning scenario, retraining from scratch can achieve the best unlearning effect, but it is expensive in both computation and storage. Existing methods try to design new unlearning mechanisms to reduce the cost from the two aspects. Although they proposed many methods, they just mitigated the influence of the main challenges and still have not solved them. Existing unlearning methods were at the beginning step that tried to implement effective unlearning. They proposed mechanisms, but most of them cannot guarantee the final results, so they also bound the unlearned item (Guo et al., 2020) or set an unlearning threshold (Nguyen et al., 2020). Although they also introduced many metrics to verify unlearning effect, the unlearning and verification processes are split. It means that the unlearning result is uncertain during the unlearning period before they finish the verification. Therefore, we find and list some open questions in centralized unlearning.
  1. (1)
    Optimizing the original three challenges. The first two challenges are hard to solve, so recent researches focus on reducing the utility degradation after unlearning. Even though many studies exist, there is still a long way to eliminate the unlearning degradation.
  2. (2)
    How to unlearn with a certain or exact goal? In other words, can we unlearn as learning? We know the purpose, to what extent we have unlearned and when we can finish or stop unlearning process.
Machine unlearning in distributed scenarios has many differences to which in centralized scenarios. We take federated unlearning as a representative example of distributed unlearning. The first difference is that federated unlearning can only be implemented locally on the client’s side if they want to unlearn some specific samples because clients’ do not upload their data to the FL server in a federated scenario. To avoid interacting with clients during unlearning, researchers (Liu et al., 2021; Wu et al., 2022b) proposed to unlearn the contribution of a whole client while not some samples of the client. The second difference is that when unlearning requests come during the FL training process, the FL server must first execute the unlearning process and broadcast the unlearned model for later updating to avoid other clients wasting computation on the before-unlearned model. The third difference is that federated unlearning is more vulnerable to catastrophic degradation than centralized unlearning because if the FL server broadcasts the catastrophic unlearned model and other clients update based on the unlearned model, it will vanish the efforts of other clients that trained before. After introducing these differences, we can see that the challenges in federated unlearning are more complex than in centralized unlearning, and we conclude the following open problems in federated unlearning.
  1. (1)
    Federated unlearning cannot use the fast retraining methods because data is out of reach for the server. Therefore, federated unlearning can only be implemented using approximate unlearning methods. However, as we know, approximate unlearning is easy to cause catastrophic unlearning, and federated learning is more vulnerable to degradation, so how to control the catastrophic in federated unlearning will be more urgent than in centralized unlearning.
  2. (2)
    Privacy in federated learning is more sensitive than in other situations because FL was designed to protect participants’ privacy at the beginning. May unlearning bring other privacy demands in federated learning, such as secret or unnoticeable unlearning?
Besides exploring machine unlearning based on regularly structured data, researchers tried to implement unlearning in graph data. Exact unlearning may be suitable for centralized graph unlearning if graph data is sparse. However, the challenges of approximate unlearning may be more difficult than structured-data-based unlearning because, in graph unlearning, the relationship and influence between data samples are more complex than structured data (Chen et al., 2021a; Chien et al., 2022). In particular, graph data includes not only the node feature value but also the connecting edge information. Therefore, in graph unlearning, the estimation of the contribution of a node will be more difficult than in regular data unlearning. The original problems in regular data unlearning will be more challenging in graph unlearning. Besides these problems, graph unlearning also face unique problems that is related to edge structure information. In graph unlearning, how to unlearn some edges or a sub-graph is a big question.
Another important part is the privacy and security issues in machine unlearning. Although machine unlearning was proposed to protect users’ privacy when they want to leave, it brings new threats in that adversaries have a chance to infer the information about the removed data. Literature (Chen et al., 2021b; Zanella-Béguelin et al., 2020) have pointed out that updates of unlearning will leak privacy information, and they proposed corresponding attacks to infer this private information. However, most recent unlearning privacy leakage attacks are similar to those in a learning situation. An attack that is tailored to unlearning mechanisms is expected. A similar situation exists in unlearning applications. Although researchers proposed to unlearn a backdoor trigger (Liu et al., 2022b) or pollution (Cao et al., 2018), they only used a few unlearning techniques and paid more effort to detect those anomalies. One important reason is that unlearning mechanism is not mature enough. We are at the beginning of machine unlearning, and there are still many problems in unlearning itself that need to be solved. Finding machine unlearning applying situation and tailoring unlearning techniques to this situation is the direction of unlearning application.
10. Summary
The survey aims to offer a comprehensive and up-to-date overview of machine unlearning techniques in a systematic way. We organize the main challenges, research advancements, corresponding techniques, and privacy and security issues in machine unlearning. Additionally, we presented a detailed and unified classification of machine unlearning. We first illustrate the complete unlearning framework, including the learning, request, unlearning, and verification. Then, we catalogue recent studies briefly into exact unlearning and approximate unlearning, and we introduce the technical details of these unlearning methods. Moreover, we noticed some new unlearning scenarios, such as federated unlearning and graph unlearning, which are also introduced in the survey. Besides, we consider privacy and security issues in machine unlearning as an important consisting part. We collect and summarize the related literature about unlearning privacy threats and applications. Ultimately, the survey provides clear summaries and comparisons between various unlearning scenarios and corresponding methods, giving a comprehensive view of current work and listing the challenges and open problems of different scenarios of machine unlearning.
We hope our survey can help classify future unlearning studies, achieve a more in-depth understanding of unlearning methods, and address complex challenges. We believe the open problems listed in Section 9 will still be challenging in the following years, and we will try to optimize some of them. Last but not least, we expect this survey can help researchers in the study of machine unlearning, regardless of the unlearning strategies or unlearning privacy and security threats or applications of unlearning.